# A High-Performance Evolutionary Multiobjective Community Detection Algorithm

### v1

@@ Abstract
Community structure is a key feature of complex networks, underpinning diverse phenomena across social, biological, and technological systems. While traditional methods like Louvain and Leiden offer efficient solutions, they rely on single-objective optimization, often failing to capture the multifaceted nature of real-world networks. Multi-objective approaches address this limitation by considering multiple structural criteria simultaneously, but their high computational cost restricts their use in large-scale settings. We propose HP-MOCD, a high-performance, fully parallel evolutionary algorithm based on NSGA-II, designed to uncover high-quality community structures by jointly optimizing conflicting objectives. HP-MOCD leverages topology-aware genetic operators and parallelism to efficiently explore the solution space and generate a diverse Pareto front of community partitions. Experimental results on large synthetic benchmarks demonstrate that HP-MOCD consistently outperforms existing multi-objective methods in runtime, while achieving superior or comparable detection accuracy. These findings position HP-MOCD as a scalable and practical solution for community detection in large, complex networks.
@@ 1 Introduction
The study of complex systems, ubiquitous in social, technological, and biological domains, heavily relies on modeling their underlying connectivity patterns as complex networks [1]. Examples range from social media platforms and the World Wide Web to protein-protein interaction networks and critical infrastructures [2]. A key characteristic of many such networks is their organization into communities, groups of nodes with dense internal connections and sparser connections to the rest of the network [3]. Identifying these communities is crucial for understanding network organization, function, and dynamics, making community detection a central research area in network science [4, 5, 6, 7]. Single-objective methods, while prevalent, often oversimplify the inherent complexity of real-world network structures, failing to capture their true multifaceted nature.
The task of community detection is commonly formulated as a combinatorial optimization problem, where the goal is to partition the network into clusters that maximize internal cohesion while minimizing interconnections between groups [8]. Since this problem is NP-hard, most practical solutions rely on traditional heuristic methods that optimize a predefined structural quality metric. In this work, we adopt a crisp partitioning model, where each node is assigned to exactly one community, reflecting a hard clustering approach that assumes mutually exclusive group membership. Among the most widely adopted are the Louvain algorithm [6], which greedily maximizes modularity ($Q$), and its improved variant, Leiden [9], which addresses issues such as disconnected communities through an enhanced refinement process.
@@
Recognizing the limitations of traditional heuristic methods, the field has progressively shifted towards multi-objective evolutionary algorithms (MOEAs) as a more expressive and flexible framework for community detection. Unlike single-metric approaches, MOEAs aim to optimize multiple, often conflicting, objectives simultaneously. This multi-criteria formulation enables a more nuanced characterization of community structures by capturing diverse structural features of the network, such as internal density, inter-community sparsity, and modularity. By design, these algorithms produce a Pareto front, i.e., a set of non-dominated solutions, offering practitioners a range of alternatives rather than a single fixed partition [12]. This enables a more comprehensive exploration of the solution space, facilitating informed decisions based on application-specific trade-offs.
@@
Early works such as MOGA-Net [13], which jointly optimized a community score and a fitness-based criterion, and Shi-MOCD [7], which explicitly targeted intra-community density and inter-community sparsity, demonstrated the potential of MOEAs to overcome structural limitations of modularity-based methods, including the resolution limit [10]. Other metaheuristics, including adaptations of the Bat Algorithm [14], further expanded the design space by exploring alternative population-based strategies. Despite their theoretical appeal, the practical adoption of MOEAs for large-scale community detection remains limited. A key challenge is their computational cost. Maintaining and evolving a population of candidate solutions requires repeated dominance comparisons, Pareto sorting, and fitness evaluations, all of which can be expensive in networks with thousands or millions of nodes and edges. Many existing implementations exhibit time complexities on the order of $O(n^{2})$ or worse, primarily due to operations involving pairwise comparisons and graph traversals [7, 15]. This makes them impractical for real-time or large-scale applications without specialized acceleration strategies.
@@
Additionally, few high-performance implementations of MOEAs are publicly available for community detection, which limits accessibility for researchers and practitioners. Most available tools are either experimental, tied to specific datasets, or lack parallelization and optimization for modern hardware. Finally, designing effective genetic operators that preserve meaningful community structures while ensuring exploration of the vast combinatorial search space remains an open problem. Balancing exploitation and exploration in this context is especially difficult due to the discrete and topologically constrained nature of the community detection problem.
@@
To address these challenges, this paper introduces the High-Performance Multi-Objective Community Detection (HP-MOCD) algorithm. HP-MOCD is a scalable evolutionary algorithm based on the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) framework [16], specifically engineered for accuracy and efficiency in large complex networks. It employs a parallel computing architecture and custom genetic operators that incorporate topological information, inspired by [7], to effectively explore the solution space. Instead of yielding a single partition, HP-MOCD produces a diverse set of Pareto-optimal solutions, each representing a different trade-off between conflicting structural objectives, thereby offering a more holistic understanding of network community structures. Moreover, this solution set can be flexibly analyzed under multiple evaluation metrics or selected according to the specific needs of a given application. Our primary contributions are as follows:
@@
    We introduce novel crossover and mutation schemes that both refine community assignments and maintain population diversity, leading to higher-quality partitions.
@@
    By embedding Pareto-front selection within the NSGA-II framework, we jointly optimize competing structural criteria, yielding a rich solution set adaptable to multiple analytical needs.
@@
    We prove that, for the sparse graphs typically encountered in practice ($|E|=O(|V|)$) with fixed population size $N_{p}$, the per-generation cost reduces to $O\bigl{(}N_{p}\,|V|\bigr{)}$, and our experiments confirm runtimes remain well below worst-case bounds.
@@
    We design a fully multi-threaded architecture that exploits modern multi-core processors, drastically reducing runtime on large-scale networks.
  5. (v)
    We provide a Free/Libre and Open Source Software (FLOSS) implementation of HP-MOCD, enabling researchers and practitioners to apply, extend, and benchmark our approach.
The remainder of this paper is structured as follows. Section 2 reviews foundational concepts and prior work in community detection, covering both traditional heuristics and multi-objective evolutionary approaches. Section 3 introduces the HP-MOCD algorithm in detail, outlining its architecture, genetic operators, and optimization strategy. Section 4 describes the experimental setup and presents a comparative evaluation against established methods. Finally, Section 5 discusses the results, highlights practical implications, and suggests directions for future research.
2 Fundamentals and Related Work
@@ 2.1 The definition of community
Formally, let $G(V,E)$ be an undirected and unweighted graph, where $V$ is the set of nodes and $E$ the set of edges. The goal of community detection is to find a partition $C=\{C_{1},C_{2},\ldots,C_{k}\}$ of the vertex set such that:
@@
In the remainder of this section, we discuss the main algorithmic approaches developed to tackle this problem. We begin with traditional heuristic methods, which have dominated the field due to their efficiency and simplicity. We then turn to more recent efforts based on multi-objective optimization, which more closely align with the direction pursued in this work.
@@ 2.2 Traditional algorithms
Another prominent strategy is label propagation, a fast heuristic where each node adopts the most frequent label among its neighbors. This process converges to a partition where nodes in the same community share the same label [22]. Label propagation is highly scalable but can yield unstable results and lacks a clear objective function, making its outcomes harder to interpret. Among modularity-based methods, the Louvain algorithm proposed by [6] has become one of the most widely used due to its balance between accuracy and computational performance. It seeks to optimize the modularity metric, defined for undirected and unweighted networks as:
In this equation, $A_{ij}$ represents the adjacency matrix of the graph, $k_{i}$ and $k_{j}$ denote the degrees of nodes $i$ and $j$, $m$ is the total number of edges, and $\delta(c_{i},c_{j})$ is an indicator function equal to 1 when nodes $i$ and $j$ belong to the same community and 0 otherwise. The Louvain method operates in two main phases. Initially, each node is placed in its own community, and the algorithm iteratively moves nodes to neighboring communities if doing so increases modularity. Once no further local improvement is possible, communities are aggregated into super-nodes, and the process is repeated on the resulting coarsened graph. This continues until convergence.
@@
Both Louvain and Leiden represent the current state of practice in scalable, modularity-based community detection. While they offer practical advantages and have become fact standard in many applications, their reliance on a single optimization criterion limits their expressiveness. In particular, they are not designed to capture multiple, potentially conflicting structural properties simultaneously. These limitations motivate the investigation of multi-objective optimization strategies, which we explore in the next section.
2.3 Multi-Objective algorithms
Traditional community detection algorithms are typically designed to optimize a single objective function, such as modularity. While efficient and widely used, these approaches exhibit critical limitations. They often produce narrow solutions that emphasize specific structural properties at the expense of others, show inconsistent performance across different network topologies, and depend heavily on the choice of objective function. These factors limit their ability to fully capture the multifaceted nature of community structure.
To overcome these shortcomings, a multi-objective optimization approach provides a more expressive and flexible framework. A multi-objective optimization problem (MOP) involves simultaneously optimizing two or more, often conflicting, objective functions. Improving one objective may lead to the deterioration of another, making it impractical to obtain a single solution that is optimal for all objectives. Instead, MOP aims to find a set of solutions that represent optimal trade-offs among the objectives. In this study, optimization is approached as the minimization of objective values.
Formally, a MOP can be stated as [24]:
$\mbox{Minimize}~{}~{}\mathbf{f}(\mathbf{x})=(f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))$
Subject to
$\mathbf{x}\in\Omega$
$PO$ =$=\{\mathbf{x}\in\Omega\mid\mathbf{f}(\mathbf{x})\in PF\}$ (5)
Multi-objective methods offer several advantages, particularly for community detection. First, they provide a diverse set of solutions, allowing users to select a partition based on context-specific priorities or further analysis. Second, the simultaneous consideration of complementary objectives can mitigate structural biases inherent in single-objective metrics and help avoid premature convergence to suboptimal solutions. Third, characteristics such as the number of detected communities can emerge naturally from the optimization process rather than being a predefined parameter. Finally, the exploration of multiple trade-off solutions often reveals richer structural insights into the network, potentially highlighting overlapping or hierarchical community structures.
Given the typically non-convex and highly combinatorial nature of the community detection problem, obtaining the true $PF$ and $PO$ is computationally challenging. Deterministic optimization methods are often impractical. Consequently, Multi-Objective Evolutionary Algorithms (MOEAs) have become a prominent strategy [25, 26, 27]. MOEAs are well-suited for exploring complex search spaces due to their population-based approach and stochastic operators. While they do not guarantee finding the exact $PF$ and $PO$ sets, MOEAs aim to produce a finite set of non-dominated solutions that effectively approximate the true Pareto front. [^1] The proximity of the set of solutions obtained by an MOEA to the true $PF$ is a measure of the algorithm’s convergence.
@@ 2.3 Multi-Objective algorithms
Building on the general definition of MOP, the Multi-Objective Community Detection (MOCD) algorithm of [7] can be formulated as follows. Let $C=\{c_{1},c_{2},\ldots,c_{k}\}$ be a partition of the node set $V$ of a graph into $k$ communities. The set of all valid partitions of $V$ is denoted as the feasible set $\Omega=\{\,C\mid C\text{ is a valid partition of }V\}$. The MOCD formulation utilizes a vector objective function $\mathbf{f}(C)$ comprising two objectives:
$f_{1}(C)$ $=1-\sum_{c\in C}\frac{\lvert E(c)\rvert}{m},\quad$ $f_{2}(C)$ $=\sum_{c\in C}\Bigl{(}\frac{\sum_{v\in c}\deg(v)}{2m}\Bigr{)}^{2}.$ (6)
In these expressions, $m=|E|$ represents the total number of edges in the network, $\deg(v)$ denotes the degree of node $v$, and $\lvert E(c)\rvert$ is the number of edges with both endpoints lying within community $c$.
The MOCD problem is then stated as:
$\min_{C\,\in\,\Omega}\;\mathbf{f}(C)\;=\;\min_{C\,\in\,\Omega}\;\bigl{(}f_{1}(C),\,f_{2}(C)\bigr{)}.$
To approximate the Pareto Optimal set, MOCD employs the Pareto Envelope-based Selection Algorithm II (PESA-II), as proposed by [28]. This evolutionary algorithm maintains both an internal population for exploration and an external archive of non-dominated solutions. Its encoding scheme is based on locus-based adjacency, which ensures that communities remain connected; however, this restricts the search space to particular types of partitions. While the use of local mutation and standard two-point crossover operators is computationally simple, it can potentially reduce genetic diversity and disrupt meaningful substructures within the solutions.
In terms of computational cost, the algorithm evaluates fitness in $O(m+n)$ time per solution, where $n=|V|$ is the number of nodes. PESA-II introduces an additional computational overhead of $O(gs^{2}(m+n))$ across $g$ generations with a population size of $s$. Furthermore, MOCD applies a Max–Min Distance selection strategy to promote diversity among solutions. This step, however, increases runtime due to the necessity of pairwise comparisons against randomly generated control solutions. Consequently, MOCD may not scale efficiently to large-scale networks.
Overall, while existing multi-objective algorithms demonstrate the theoretical advantages of Pareto optimization, they remain under-explored in terms of implementation quality, scalability, and general usability. This highlights the need for more efficient and accessible methods, particularly in large-scale applications, a gap that this work aims to address.

### v2

@@ Abstract
Community detection in complex networks is fundamental across social, biological, and technological domains. While traditional single-objective methods like Louvain and Leiden are computationally efficient, they suffer from resolution bias and structural degeneracy. Multi-objective evolutionary algorithms (MOEAs) address these limitations by simultaneously optimizing conflicting structural criteria, however, their high computational costs have historically limited their application to small networks. We present HP-MOCD, a High-Performance Evolutionary Multiobjective Community Detection Algorithm built on Non-dominated Sorting Genetic Algorithm II (NSGA-II), which overcomes these barriers through topology-aware genetic operators, full parallelization, and bit-level optimizations—achieving theoretical $O(G\cdot N_{p}|V|)$ complexity. We conduct experiments on both synthetic and real-world networks. Results demonstrate strong scalability, with HP-MOCD processing networks of over 1,000,000 nodes while maintaining high quality across varying noise levels. It outperforms other MOEAs by more than 531 times in runtime on synthetic datasets, achieving runtimes as low as 57 seconds for graphs with 40,000 nodes on moderately powered hardware. Across 14 real-world networks, HP-MOCD was the only MOEA capable of processing the six largest datasets within a reasonable time, with results competitive with single-objective approaches. Unlike single-solution methods, HP-MOCD produces a Pareto Front, enabling individual-specific trade-offs and providing decision-makers with a spectrum of high-quality community structures. It introduces the first open-source Python MOEA library compatible with networkx and igraph for large-scale community detection.
@@ 1 Introduction
The study of complex systems, ubiquitous in social, technological, and biological domains, heavily relies on modeling their underlying connectivity patterns as complex networks [1]. Examples range from social media platforms and the World Wide Web to protein-protein interaction networks and critical infrastructures [2]. A key characteristic of many such networks is their organization into communities, groups of nodes with dense internal connections and sparser connections to the rest of the network [3]. Figure 1 provides a canonical example of community structure.
Figure 1: A conceptual illustration of community structure. The network is partitioned into two distinct communities (blue and green). The solid gray lines represent dense intra-community edges, which create cohesive groups. The single dashed red line represents a sparse inter-community edge (or bridge), highlighting the weak connectivity between the groups. The goal of community detection is to identify such partitions.
Identifying these communities is crucial for understanding network organization, function, and dynamics, making community detection a central research area in network science [4, 5, 6, 7]. The task is commonly formulated as a combinatorial optimization problem, where the goal is to partition the network into clusters that maximize internal cohesion while minimizing interconnections between groups [8]. Since this problem is NP-hard, most practical solutions rely on heuristic methods that optimize a predefined structural quality metric [9].
Among these, single-objective approaches are particularly widespread due to their efficiency and simplicity. However, they often oversimplify the multifaceted nature of real-world networks by focusing on a single structural criterion. Two of the most widely used single-objective methods are the Louvain algorithm [6], which greedily maximizes modularity, and its improved variant, Leiden [10], which addresses issues such as disconnected communities through an enhanced refinement process.
@@
Recognizing the limitations of traditional heuristic methods, the field has progressively shifted towards multi-objective evolutionary algorithms (MOEAs) as a more expressive and flexible framework for community detection [13, 14]. Unlike single-metric approaches, MOEAs aim to optimize multiple, often conflicting, objectives simultaneously. This multi-criteria formulation enables a more nuanced characterization of community structures by capturing diverse structural features of the network, such as internal density, inter-community sparsity, and modularity. By design, these algorithms produce a Pareto front, i.e., a set of non-dominated solutions, offering practitioners a range of alternatives rather than a single fixed partition [15]. This enables a more comprehensive exploration of the solution space, facilitating informed decisions based on application-specific trade-offs.
@@
Early works such as MOGA-Net [16], which jointly optimized a community score and a fitness-based criterion, and Shi-MOCD [7], which explicitly targeted intra-community density and inter-community sparsity, demonstrated the potential of MOEAs to overcome structural limitations of modularity-based methods, including the resolution limit [11]. Following these foundational works, the field has seen a range of proposals aimed at improving search efficiency and solution quality. In [17], the authors developed learning-based strategies to select nodes in each round of the multi-objective evolutionary process, called Local Search for Community Detection (LSCD). The work [18] introduces a continuous-encoding-based approach with a clustering coefficient in a multi-objective framework, called CECC-Net, which represents the community structure. Focusing on the challenge of detecting overlapping communities, the work [19] presents an MOEA that utilizes an ensemble of different initializations. Despite their theoretical appeal, the practical adoption of those MOEAs for large-scale community detection remains limited, with applications to synthetic and real-world networks of no more than 10,000 nodes. A key challenge is their computational cost. Maintaining and evolving a population of candidate solutions requires repeated dominance comparisons, Pareto sorting, and fitness evaluations, all of which can be expensive in networks with thousands or millions of nodes and edges. Many existing implementations exhibit time complexities on the order of $O(n^{2})$ or worse, primarily due to operations involving pairwise comparisons and graph traversals [7, 20]. This makes them impractical for real-time or large-scale applications without specialized acceleration strategies.
More recent algorithms include KRM and CCM [13], both based on the NSGA-III framework. KRM optimizes three objectives: Kernel-k-means, Ratio-cut, and Modularity using a reference-point guided selection mechanism, while CCM targets Community fitness, Community score, and Modularity. Another notable contribution is CDRME [14], a four-phase evolutionary method that incorporates similarity-based initialization and random-walk-based seeding to balance cut density, ratio cut, and modularity across network scales.
Despite their theoretical appeal, the practical adoption of these MOEAs for large-scale community detection remains limited, with applications to synthetic and real-world networks of no more than 10,000 nodes. A key challenge is their computational cost. Maintaining and evolving a population of candidate solutions requires repeated dominance comparisons, Pareto sorting, and fitness evaluations, all of which can be expensive in networks with thousands or millions of nodes and edges. Many existing implementations exhibit time complexities on the order of $O(n^{2})$ or worse, primarily due to operations involving pairwise comparisons and graph traversals [7, 20]. This makes them impractical for real-time or large-scale applications without specialized acceleration strategies.
@@
Additionally, few implementations of MOEAs are publicly available for community detection, which limits accessibility for researchers and practitioners. Most available tools are either experimental, tied to specific datasets, or lack parallelization and optimization for modern hardware. Finally, designing effective genetic operators that preserve meaningful community structures while ensuring exploration of the vast combinatorial search space remains an open problem. Balancing exploitation and exploration in this context is especially difficult due to the discrete and topologically constrained nature of the community detection problem.
@@
To address these challenges, this paper introduces the High-Performance Multi-Objective Community Detection (HP-MOCD) algorithm. HP-MOCD is a scalable evolutionary algorithm based on the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) framework [21], specifically engineered for accuracy and efficiency in large complex networks. It employs a parallel computing architecture and custom genetic operators that incorporate topological information to explore the solution space effectively. Instead of yielding a single partition, HP-MOCD produces a diverse set of Pareto-optimal solutions, each representing a different trade-off between conflicting structural objectives, thereby offering a more holistic understanding of network community structures. Moreover, this solution set can be flexibly analyzed under multiple evaluation metrics or selected according to the specific needs of a given application. Our primary contributions are as follows:
@@
    We introduce a novel topology-aware crossover and mutation schemes that both refine community assignments and preserve population diversity, resulting in more accurate and structurally consistent community partitions.
@@
    We prove that, for the sparse graphs typically encountered in practice ($|E|=O(|V|)$) with fixed population size $N_{p}$, the per-generation cost reduces to $O\bigl{(}N_{p}\,|V|\bigr{)}$. Our experiments confirm runtimes remain well below worst-case bounds.
@@
    We design a fully multi-threaded NSGA-II architecture that exploits modern multi-core processors, drastically reducing runtime on large-scale networks, .
@@
    We provide a Free/Libre and Open-Source Software (FLOSS) implementation of HP-MOCD, enabling researchers and practitioners to apply, extend, and benchmark our approach.
The remainder of this paper is structured as follows. Section 2 reviews foundational concepts in community detection, including definitions and key modeling assumptions. Section 3 surveys prior work, covering both traditional algorithms and recent multi-objective approaches. Section 4 formalizes the multi-objective problem and presents the HP-MOCD algorithm in detail, including its architecture, genetic operators, and parallel execution model. Section 5 describes the experimental setup and reports a comparative evaluation against several baselines on synthetic and real-world networks. Finally, Section 6 discusses the implications, limitations, and potential directions for future research.
2 Background
@@ 2.1 The Definition of Community
Formally, let $G=(V,E)$ be an undirected and unweighted graph, where $V$ is a finite, non-empty set of nodes and $E\subseteq\{\{u,v\}\mid u,v\in V,u\neq v\}$ is the set of undirected edges.
@@
2.2 Multi-Objective Optimization
A multi-objective optimization problem (MOP) involves simultaneously optimizing two or more, often conflicting, objective functions. Improving one objective may lead to the deterioration of another, making it impractical to obtain a single solution that is optimal for all objectives. Instead, MOP aims to find a set of solutions that represent optimal trade-offs among the objectives. In this study, optimization is approached as the minimization of objective values.
Formally, a MOP can be stated as [25]:
$\mbox{Minimize}~~\mathbf{f}(\mathbf{x})=(f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))\mbox{ subject to }\mathbf{x}\in\Omega$ (2)
$PO$ =$=\{\mathbf{x}\in\Omega\mid\mathbf{f}(\mathbf{x})\in PF\}$ (4)
3 Related Work
Community detection is a vast field. Here, we compare and contrast HP-MOCD with the two main classes of algorithms relevant to our work: traditional algorithms and other MOEAs.
@@ 3.1 Traditional Algorithms
Another prominent strategy is label propagation, a fast heuristic where each node adopts the most frequent label among its neighbors. This process converges to a partition where nodes in the same community share the same label [27]. Label propagation is highly scalable but can yield unstable results and lacks a clear objective function, making its outcomes harder to interpret. Among modularity-based methods, the Louvain algorithm proposed by [6] has become one of the most widely used due to its balance between accuracy and computational performance. It seeks to optimize the modularity metric [28], which measures the quality of a partition by comparing the density of edges within communities to the expected density in a randomized null model that preserves the node degrees. In essence, modularity quantifies how much more densely connected the nodes are within communities than would be expected by chance. The modularity score theoretically ranges from $-\frac{1}{2}$ to 1, where higher values indicate a stronger community structure, and values near zero or negative suggest that the detected communities are no better—or even worse—than random.
@@
Algorithms such as Louvain and Leiden represent the current state of practice in scalable, modularity-based community detection. Their combination of simplicity, computational efficiency, and availability in widely used libraries has made them widely adopted solutions in many real-world applications. Despite this practical success, their reliance on a single optimization criterion inherently limits their expressiveness. In particular, they are not designed to capture multiple, potentially conflicting structural properties simultaneously, nor do they offer the user flexibility in selecting solutions that trade off different qualities such as compactness, separation, or size balance. These methods produce only a single partition per execution, often overlooking alternative high-quality solutions that might better suit specific applications or structural goals. These limitations motivate the investigation of multi-objective optimization strategies, which aim to overcome the rigidity of single-objective heuristics by generating diverse sets of non-dominated solutions. In the following section, we review some of the recent advances in this area and highlight their potential to provide richer structural insights and greater flexibility for real-world community detection.
3.2 Multi-Objective Evolutionary Algorithms
The limitations of single-objective methods motivated the development of Multi-Objective Evolutionary Algorithms (MOEAs) for community detection. Unlike single-metric approaches, MOEAs optimize multiple, often conflicting, objectives simultaneously, such as internal density and external sparsity. This approach yields a Pareto front [^1] The proximity of the set of solutions obtained by an MOEA to the true $PF$ is a measure of the algorithm’s convergence., a set of non-dominated solutions—offering a range of robust alternatives rather than a single, potentially biased partition [15, 30]. The use of MOEAs has become a prominent strategy due to their population-based nature, which is well-suited for exploring the complex and combinatorial solution space of community detection [31, 32].
@@ 3.2 Multi-Objective Evolutionary Algorithms
The Multi-Objective Community Detection (MOCD) algorithm by [7] jointly minimizes two objectives related to intra-community density and inter-community sparsity. MOCD employs the PESA-II evolutionary framework and locus-based operators. However, its selection strategy and operators lead to high computational overhead, with a complexity of $O(gs^{2}(m+n))$, making it impractical for large networks.
More recent research has continued to explore the potential of MOEAs, introducing more sophisticated objective functions and hybrid strategies. The work by [13], introduced a customized NSGA-III framework to simultaneously optimize three objectives (e.g., Kernel k-means, Ratio-cut, and Modularity). Their contribution demonstrated the feasibility of using many-objective optimization to find community structures that represent a more nuanced trade-off between different structural properties. [14] proposed a hybrid method that first uses random walks to generate an initial set of high-quality partitions, which are then refined by a multi-objective evolutionary algorithm. This approach aims to accelerate convergence by seeding the evolutionary process in a more promising region of the search space.